how long will it take for a wooden block of mass 2kg to from rest a distance of 3m down a rough plane inclined at 30 degree to the horizontal?given coefficient of sliding friction=0.20

Grade:11

1 Answers

Anvita Bajpai
6 years ago
u=0, s=3
applying s=ut+1/2at2
since u=0, the formula reduces to,
s=1/2 at2  where a=gsin 30.-umgcos 30
t=(2s/a)1/2